Fermi National Laboratory
 

Accelerator Update for May 1, 2009 - May 4, 2009

 

The Integrated Luminosity 4/27/09 to 5/04/09 was 67.28 pb-1.

Friday May 1

The day shift began with Operations monitoring the Tevatron (TeV) store 7009 (212.96E30) and the Recycler stash (94.5E10), with the Antiproton Source stacking (25.1E10), with the Linac, Booster, and Main Injector (MI) operating normally, with the Meson MTest, MiniBooNE, and NuMI experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams.

Operations terminated store 7009 at 4:04 PM.

At 4:54 PM, vacuum techs began a 4 hours access into Tevatron enclosures A-E for vacuum system work, CHL also began wet engine work at B4. CDF and D0 used this opportunity to make controlled accesses.

CHL wet engine repair was completed by 7:04 PM.

Tevatron A-E access was completed by 9:01 PM.

Saturday May 2

The midnight shift began with the TeV preparing for shot setup, with Operations monitoring the stash (429.3E10) and the stack (25.5E10), and with NuMI and MiniBooNE taking beam.

Operations had put the TeV into HEP shot set-up by 2:45 AM.

Operations established store 7013 at 3:35 AM with an initial luminosity of 325.83E30.

MTest turn on was delayed due to problems with the Meson Frontend, MTest began taking beam at 7:00 AM.

Operations terminated store 7013 at 3:48 PM.

Operations put the TeV into HEP shot set-up by 4:00 PM.

Operations established store 7014 at 5:21 PM with an initial luminosity of 320.5E30.

Operations delivered beam to Pbar, NuMI, and MiniBooNE until the Klystron Buncher failed at 11:33 PM.

Sunday May 3

The midnight shift began with Operations monitoring store 7014, and the stash, and the stack. Linac was down due to the Klystron Buncher problem and with NuMI and MiniBooNE off due to the Klystron problems.

Experts repaired the Klystron Buncher and beam returned at 2:40 AM.

Operations terminated store 7014 at 9:36 AM.

DZero accessed their collider hall at 10:40 AM.

DZero access was completed at 11:32 AM.

Operations put the TeV into HEP shot set-up by 1:40 PM.

Operations established store 7018 at 3:06 PM with an initial luminosity of 312.2E30.

Monday May 4

The midnight shift began with Operations monitoring the store (7018), and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam.

Operations terminated store 7018 at 4:40 AM.

Operations put the TeV into HEP shot set-up by 4:51 AM.

Operations established store 7019 at 6:13 AM with an initial luminosity of 304.86E30.

The Plans for Monday

The plans for today are to stack, stash, and store.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 05/04/2009    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory